Log in
wiki.ucar.edu
Spaces
Hit enter to search
Help
Online Help
Keyboard Shortcuts
Feed Builder
What’s new
About Confluence
Log in
SOAS Field Project - ISFS
Edit space details
Pages
Blog
Blog
2013
July
Soil Installations
SOAS post-project sonic temperature chamber tests
Tower Mote - Wand Serial Numbers and Locations
DONE
Tower Down Base Gone
Base Packed
Tower Removal
Tower Undressed
SHUTDOWN!
Tower Climbed
Tower Net and Power outage - no data loss
SOAS ISFS Status 15July13
Tower Ride OPEN
20m IRT outage
SOAS ISFS Status 14July13
TRAM Trailer Tire
Soilb mote rebooted
20m IRT removed
SOAS ISFS Status 13July13
Tower work - Kh20 cleaning. licor, par wetness partial
SOAS ISFS Status 12July13
SOAS ISFS Status 11July13
20m Sonic Electronics Box Adjustment
Tower Work Wed. 10July13 1500-1600
SOAS ISFS Status 10July13
Mote Problems - outage of Soil-A
SOAS ISFS Status 9July13
For Chris' Post Project Checking
43m barometer has undecipherable char messages
Tower Work
On Da Tower
26m TRH outage - moved to port 11
SOAS ISFS Status 8July13
Rad Mote outage
20m Sonic Down
SOAS ISFS Status July7
SOAS ISFS Status July06
Tower visit
SOAS ISFS Status July05
tower climbing
lost power last night
SOAS ISFS Status July 04
second rain event this afternoon
updated netcdf files, restarted processes
rain event
Tsfc.8m down
tower work
SOAS ISFS Status July 3
SOAS ISFS Status July 02
Duff Tsoil Probe
TP01 question
SOAS ISFS Status July 1
June
May
April
Browse pages
Configure
Space tools
A
t
tachments (0)
Page History
Page Information
Resolved comments
View Source
Export to PDF
Hide Inline Comments
Blog
2013
July
20
Jira links
DONE
Created by
Chris Golubieski
, last updated on
Jul 20, 2013
1 minute read
July 20, 2013
We are all packed and site is cleaner than when we left it. Good bye, SOAS.
No labels
Overview
Content Tools
{"serverDuration": 69, "requestCorrelationId": "704335bc42b09bd1"}